A marble is drawn at random from a box containing 10 red,30 white,20 blue& 15 orange marbles .find the probability that it is 1)orange or red 2)not red or blue

1
Expert's answer
2021-12-30T17:40:04-0500

There are "10+30+20+15=75" marbles.

1)

"P(orange\ or\ red)=\dfrac{15+10}{75}=\dfrac{1}{3}"

2)

"P(not(\ red\ or\ blue))=P(orange\ or\ white)"

"=\dfrac{15+30}{75}=\dfrac{3}{5}"